# Transformer Copilot: Learning from The Mistake Log in LLM Fine-tuning

### v1

@@ Abstract
Large language models are typically adapted to downstream tasks through supervised fine-tuning on domain-specific data. While standard fine-tuning focuses on minimizing generation loss to optimize model parameters, we take a deeper step by retaining and leveraging the model’s own learning signals, analogous to how human learners reflect on past mistakes to improve future performance. We first introduce the concept of Mistake Log to systematically track the model’s learning behavior and recurring errors throughout fine-tuning. Treating the original transformer-based model as the Pilot, we correspondingly design a Copilot model to refine the Pilot’s inference performance via logits rectification. We name the overall Pilot-Copilot framework the Transformer Copilot, which introduces (i) a novel Copilot model design, (ii) a joint training paradigm where the Copilot continuously learns from the evolving Mistake Log alongside the Pilot, and (iii) a fused inference paradigm where the Copilot rectifies the Pilot’s logits for enhanced generation. We provide both theoretical and empirical analyses on our new learning framework. Experiments on 12 benchmarks spanning commonsense, arithmetic, and recommendation tasks demonstrate that Transformer Copilot consistently improves performance by up to 34.5%, while introducing marginal computational overhead to Pilot models and exhibiting strong scalability and transferability [^1] Code will be released at https://github.com/jiaruzouu/TransformerCopilot.
@@ 1 Introduction
Transformers, the foundation of modern large language models (LLMs), leverage attention and feedforward layers to compute logits for sequence generation vaswani2017attention . Pre-trained on general-domain corpora, these models capture rich statistical patterns and exhibit strong generation capabilities brown2020language ; wei2021finetuned ; huang2022towards ; mo2024large . On top of that, supervised fine-tuning (SFT) serves as a critical technique for adapting pre-trained LLMs to specific domains kenton2019bert ; raffel2020exploring ; wei2021finetuned ; zou2024stempom . While SFT enables significant flexibility and task-specific optimization, the performance of fine-tuned LLMs during inference often remains suboptimal, exhibiting misalignment between training and testing stages li2024hindsight ; wang2024loss . This gap arises from the model’s inability to fully capture task-specific nuances or from overfitting to patterns within the training data, ultimately degrading its final performance raffel2020exploring ; min2022rethinking ; zhao2021calibrate ; mccoy2019right . Without data-side interventions madaan2023self ; miao2023selfcheck ; gulcehre2023reinforced or receiving external feedback ouyang2022training ; shinn2023reflexion ; yao2023react , this paper aims to address a fundamental question: Can we enhance the inference performance by retaining and leveraging the model’s own learning signals in standard fine-tuning?
@@ 3.1 The Copilot Model Design
New $\text{New }Q$ $=H^{C}_{l-1}\cdot W^{Q},\text{ for }l=1,...,L^{C},$ (6)
New $\text{New }K$ $=\text{Concat}\big{(}\widetilde{X}_{t},\ \text{Pool}_{L^{P}}\big{(}h_{t}(\widetilde{X}_{t};\theta^{P}_{t-1})\big{)}\big{)}\cdot W^{K},$
New $\text{New }V$ $=\text{Concat}\big{(}\widetilde{X}_{t},\ \text{Pool}_{L^{P}}\big{(}h_{t}(\widetilde{X}_{t};\theta^{P}_{t-1})\big{)}\big{)}\cdot W^{V},$
@@ 4       $\triangledown$ Pilot - token-level forward pass
8      $\triangledown$ Collect Mistake Log (§$\S$2.2)
@@
13      /* For brevity, we reuse notation $t$/
@@
15       $\triangledown$ Copilot - learn from the Mistake Log (§$\S$3.1)
@@ 17              $f^{C}_{t,i}\leftarrow f^{C}(\widetilde{X}_{t},h_{t,<i},\ell_{t,<i};\theta_{t-1}^{C})$
Input: $\theta^{P}_{T}$, $\theta^{C}_{T}$; Tuning parameter $\lambda$ 1 Draw new $X_{t}\sim\mathcal{D}_{\mathcal{X}},t>T$ 2 for  $i=1,\dots,n$ do 3       $\hat{p}_{t,i}\leftarrow\text{softmax}(f^{P}(X_{t},\hat{y}_{t,<i};\theta^{P}_{T}))$ 4       Observe $\widetilde{X}_{t},h_{t,<i}$ from $f^{P}$ 5       $f^{C}_{t,i}\leftarrow f^{C}(\widetilde{X}_{t},h_{t,<i},f^{C}_{t,<i};\theta^{C}_{T})$ 6       $\tilde{p}_{t,i}\leftarrow\hat{p}_{t,i}+\lambda f^{C}_{t,i}$ (via Eq.8) 7       $\hat{y}_{t,i}\leftarrow$ Decoding$(\tilde{p}_{t,i})$ 8 end for return $(\hat{y}_{t,1},\dots,\hat{y}_{t,n})$ Algorithm 2 Inference Paradigm
@@ 4 Analyses - Why Learn from the Mistake Log?
Theorem 4.1 suggests that the rectified prediction $\tilde{p}_{t,i}$ after incorporating the Copilot model achieves strictly lower expected error at $k$-dimension under mild assumptions and a proper $\lambda$, indicating the Copilot helps improve the inference performance of the Pilot. The full proof is provided in Appendix C. In addition, Remark 4.2 implies that the Copilot model can improve inference performance without needing to match the Pilot’s accuracy in isolation. This insight motivates us to apply a relatively smaller size of a Copilot to complement the Pilot in our empirical implementation.
Figure 3: Logits Correction by Copilot. We visualize the logits correction introduced by a 1B Copilot model (computed as ||||Fused logits -- Pilot logits||||) to highlight the shift by the Copilot’s rectification. Left: Percentage of logits correction over original Pilot’s output logits range for three LLaMA-3 Pilot models. Right: Distribution of logits correction magnitudes across reasoning types.
@@
Empirical Analysis. Complementing our theoretical analysis, we empirically examine the rectification effectiveness of the Copilot model during inference. We leave the setups in Appendix D. Figure 3 illustrates the average logits correction induced by the 1B Copilot model across different Pilot models and reasoning categories. Given that the typical logits range is approximately $[-10,10]$, the observed shifts on the logits distribution indicate a clear and consistent adjustment on the final predictions by the Copilot model.
We further verify that this Copilot’s adjustment indeed steers the token prediction toward the correct direction: We analyze representative error patterns frequently observed in the Pilot model’s output, particularly factual and formatting mistakes. Figure 4 shows a detailed example of token-level logits rectification on Pilot model LLaMA-3.2-3B by the 1B Copilot, visualized using the layer-wise Logits Lens belrose2023eliciting . At mid-inference, the Pilot does not follow the correct answer format and makes mistakes (the correct token ‘answer’ has a high but suboptimal logit). The Copilot rectifies the prediction by decreasing the logit of the incorrect token ‘forgot’ and amplifying that of the correct token, thereby correcting the token prediction error. We leave analyses on other error patterns in Appendix D.
@@ 5 Empirical Evaluations
Tasks and Datasets. To comprehensively evaluate T-Copilot, we utilize a broad suite of reasoning and generation tasks: (i) Commonsense reasoning: PIQA bisk2020piqa , HellaSwag zellers2019hellaswag , WinoGrande sakaguchi2021winogrande , BoolQ clark2019boolq , SIQA sap2019socialiqa , and OpenbookQA (OBQA) mihaylov2018can . (ii) Arithmetic reasoning: AQuA ling2017program , GSM8K cobbe2021training , MAWPS koncel2016mawps , and SVAMP patel2021nlp . and (iii) Downstream Recommendation: Beauty he2016ups and LastFM rozemberczki2020characteristic . Detailed dataset descriptions are provided in Appendix E.
Implementation Details. For T-Copilot, we construct the Copilot model using the same type of decoder architecture as the Pilot model to ensure consistency. We use the AdamW optimizer and Cosine learning rate scheduler for both Pilot and Copilot models. We modify the generate in HuggingFace Transformers huggingface-transformers to perform token-level logits fusion and rectified next-token generation during inference. All experiments are conducted on NVIDIA A100 GPUs. We leave all hyperparameter setups and training/inference details in Appendix F.1.
@@
Models and Baselines. We incorporate T-Copilot with varying backbone Pilot models. For encoder-decoder Pilots, we utilize T5 raffel2020exploring and FLAN-T5 chung2024scaling across small/base/large variants. For decoder-only Pilots, we employ multiple models from LLaMA-3 dubey2024llama and Qwen2.5 yang2024qwen2 families. We denote T-Copilot-small/base/0.5B/1B/3B as the Copilot model on different scales. Detailed model configuration and implementation details are provided in Appendix F.2. We compare against three baseline types: (i) Pilot-only models as described above. (ii) Frontier LLMs with comparable and larger parameters, including LLaMA-3.1-8B dubey2024llama , Gemma-2-9B team2024gemma , and Qwen2.5-14B. (iii) Layer/Adapter expansion methods, including MoE models shazeer2017outrageously (Mistral-7B, Ministral-8B), LLaMA/Mistral-Pro-8B wu2024llama , Mergekit-9B goddard2024arcee , and TIESyadav2024ties . Detailed baseline descriptions are provided in Appendix F.3.
@@ 5.1 Incorporating T-Copilot into Pilot Models Yields Better Performance
Comparison with Size-Matched Baselines. As shown in Table 2, we first compare our method against stronger models with larger parameters under the same model backbones. While LLaMA-3.2-3B initially lags significantly behind LLaMA-3.1-8B, incorporating T-Copilot-3B enables the model to outperform LLaMA-3.1-8B, despite using 2B fewer total parameters. Similarly, for the Qwen2.5 series, incorporating T-Copilot-3B enables the smaller Qwen2.5-7B to surpass Qwen2.5-14B with 4B fewer parameters. To provide a broader perspective, we also compare with strong baselines from different methods and model types. For instance, although LLaMA-3.2-3B originally trails behind models like Ministral-8B and LLaMA-Pro-8B, incorporating T-Copilot-3B enables it to outperform the strongest baseline under the 8B scale, Ministral-8B, while maintaining a 2B parameter advantage. Due to page limits, full comparison results are provided in Appendix G.1.
Downstream Tasks. Additional evaluation of T-Copilot and baseline comparisons on downstream recommendation tasks is provided in Appendix G.2.
@@ 5.2 Efficiency, Transferability, and Scalability
Efficiency. To thoroughly evaluate T-Copilot’s running efficiency, we compare against Pilot and baseline models with the same LLaMA-3 backbone architecture under similar parameter scales. As shown in Figure 5, T-Copilot maintains comparable inference throughput (Figure 5 (a)) and training speed (Figure 5 (b)) to its corresponding Pilot models, while incurring only a 4% marginal average increase in time overhead (Figure 5 (c)). In contrast, other baselines such as LLaMA-Pro-8B and MergeKit-9B suffer from significantly higher latency and computational costs relative to their base model LLaMA-3.1-8B. We provide a more detailed inference latency report in Appendix G.5 (Table 15) and discuss the architectural advantage of our model design in Appendix B.1.
Transferability and Scalability. Due to space constraints, we move all experiment details here in Appendix G.3 and Appendix G.4. Our results show that T-Copilot is scalable and can be seamlessly transferred to new Pilot models with comparable effectiveness, with no additional fine-tuning needed. Ablation Studies. Detailed ablation studies on T-Copilot, including model design choices, input insertion patterns, and the effect of the hyperparameter $\lambda$, are presented in Appendix G.6.
@@ 6 Related Works
Recent research has explored various self-refinement techniques in LLMs to generate high-quality outputs. Models either iteratively prompt themselves with updated responses madaan2023self ; gulcehre2023reinforced ; shinn2023reflexion or optimize their behavior using external human or synthetic feedback ouyang2022training ; yao2023react ; miao2023selfcheck . Orthogonal to external supervision such as additional prompting, multi-stage feedback, or explicit reward optimization, our work focuses on capturing model-internal signals during fine-tuning to achieve token-level rectification, without modifying the training objective or data distribution. We leave the additional related work and discussions in the Appendix H.
@@ 7 Conclusion
In this paper, we introduce Transformer Copilot, a novel learning framework that enhances Transformer-based Pilot models by integrating an auxiliary Copilot model during fine-tuning. By capturing the Pilot model’s learning signals in a Mistake Log during fine-tuning, the Copilot model learns to rectify the Pilot’s logits at inference time, enabling error-aware predictions. We provide both theoretical and empirical evidence that our method improves the Pilot model’s inference predictions. Experiments on 12 benchmarks demonstrate the effectiveness, efficiency, scalability, and transferability of Transformer Copilot. Discussions on limitations are provided in Appendix A.

### v2

@@ Abstract
Large language models are typically adapted to downstream tasks through supervised fine-tuning on domain-specific data. While standard fine-tuning focuses on minimizing generation loss to optimize model parameters, we take a deeper step by retaining and leveraging the model’s own learning signals, analogous to how human learners reflect on past mistakes to improve future performance. We first introduce the concept of Mistake Log to systematically track the model’s learning behavior and recurring errors throughout fine-tuning. Treating the original transformer-based model as the Pilot, we correspondingly design a Copilot model to refine the Pilot’s inference performance via logits rectification. We name the overall Pilot-Copilot framework the Transformer Copilot, which introduces (i) a novel Copilot model design, (ii) a joint training paradigm where the Copilot continuously learns from the evolving Mistake Log alongside the Pilot, and (iii) a fused inference paradigm where the Copilot rectifies the Pilot’s logits for enhanced generation. We provide both theoretical and empirical analyses on our new learning framework. Experiments on 12 benchmarks spanning commonsense, arithmetic, and recommendation tasks demonstrate that Transformer Copilot consistently improves performance by up to 34.5%, while introducing marginal computational overhead to Pilot models and exhibiting strong scalability and transferability.
@@ 1 Introduction
Transformers, the foundation of modern large language models (LLMs), leverage attention and feedforward layers to compute logits for sequence generation vaswani2017attention . Pre-trained on general-domain corpora, these models capture rich statistical patterns and exhibit strong generation capabilities brown2020language ; wei2021finetuned ; mo2024large . On top of that, supervised fine-tuning (SFT) serves as a critical technique for adapting pre-trained LLMs to specific domains kenton2019bert ; raffel2020exploring ; wei2021finetuned . While SFT enables significant flexibility and task-specific optimization, the performance of fine-tuned LLMs during inference often remains suboptimal, exhibiting misalignment between training and testing stages li2024hindsight ; wang2024loss . This gap arises from the model’s inability to fully capture task-specific nuances or from overfitting to patterns within the training data, ultimately degrading its final performance raffel2020exploring ; min2022rethinking ; zhao2021calibrate ; mccoy2019right . Without data-side interventions madaan2023self ; miao2023selfcheck ; gulcehre2023reinforced or receiving external feedback ouyang2022training ; shinn2023reflexion ; yao2023react , this paper aims to address a fundamental question: Can we enhance the inference performance by retaining and leveraging the model’s own learning signals in standard fine-tuning?
@@ 3.1 The Copilot Model Design
$\text{New }Q$ $=H^{C}_{l-1}\cdot W^{Q},\text{ for }l=1,.,L^{C},$ (6)
$\text{New }K$ $=\text{Concat}\big(\widetilde{X}_{t},\ \text{Pool}_{L^{P}}\big(h_{t}(\widetilde{X}_{t};\theta^{P}_{t-1})\big)\big)\cdot W^{K},$
$\text{New }V$ $=\text{Concat}\big(\widetilde{X}_{t},\ \text{Pool}_{L^{P}}\big(h_{t}(\widetilde{X}_{t};\theta^{P}_{t-1})\big)\big)\cdot W^{V},$
@@ 4  $\triangledown$ Pilot - token-level forward pass
8 $\triangledown$ Collect Mistake Log ($\S$2.2)
@@
13 /* For brevity, we reuse notation $t$ */
@@
15  $\triangledown$ Copilot - learn from the Mistake Log ($\S$3.1)
@@ 17     $f^{C}_{t,i}\leftarrow f^{C}(\widetilde{X}_{t},h_{t,<i},\ell_{t,<i};\theta_{t-1}^{C})$
[…]
@@ 4 Analyses - Why Learn from the Mistake Log?
Theorem 4.1 suggests that the rectified prediction $\tilde{p}_{t,i}$ after incorporating the Copilot model achieves strictly lower expected error at $k$-dimension under mild assumptions and a proper $\lambda$, indicating the Copilot helps improve the inference performance of the Pilot. The full proof is provided in Appendix B. In addition, Remark 4.2 implies that the Copilot model can improve inference performance without needing to match the Pilot’s accuracy in isolation. This insight motivates us to apply a relatively smaller size of a Copilot to complement the Pilot in our empirical implementation.
Figure 3: Logits Correction by Copilot. We visualize the logits correction introduced by a 1B Copilot model (computed as ||Fused logits - Pilot logits||) to highlight the shift by the Copilot’s rectification. Left: Percentage of logits correction over original Pilot’s output logits range for three LLaMA-3 Pilot models. Right: Distribution of logits correction magnitudes across reasoning types.
@@
Empirical Analysis. Complementing our theoretical analysis, we empirically examine the rectification effectiveness of the Copilot model during inference. We leave the setups in Appendix C. Figure 3 illustrates the average logits correction induced by the 1B Copilot model across different Pilot models and reasoning categories. Given that the typical logits range is approximately [10,10][-10,10], the observed shifts on the logits distribution indicate a clear and consistent adjustment on the final predictions by the Copilot model.
We further verify that this Copilot’s adjustment indeed steers the token prediction toward the correct direction: We analyze representative error patterns frequently observed in the Pilot model’s output, particularly factual and formatting mistakes. Figure 4 shows a detailed example of token-level logits rectification on Pilot model LLaMA-3.2-3B by the 1B Copilot, visualized using the layer-wise Logits Lens belrose2023eliciting . At mid-inference, the Pilot does not follow the correct answer format and makes mistakes (the correct token ‘answer’ has a high but suboptimal logit). The Copilot rectifies the prediction by decreasing the logit of the incorrect token ‘forgot’ and amplifying that of the correct token, thereby correcting the token prediction error. We leave analyses on other error patterns in Appendix C.
@@ 5 Empirical Evaluations
Tasks and Datasets. To comprehensively evaluate T-Copilot, we utilize a broad suite of reasoning and generation tasks: (i) Commonsense reasoning: PIQA bisk2020piqa , HellaSwag zellers2019hellaswag , WinoGrande sakaguchi2021winogrande , BoolQ clark2019boolq , SIQA sap2019socialiqa , and OpenbookQA (OBQA) mihaylov2018can . (ii) Arithmetic reasoning: AQuA ling2017program , GSM8K cobbe2021training , MAWPS koncel2016mawps , and SVAMP patel2021nlp . and (iii) Downstream Recommendation: Beauty he2016ups and LastFM rozemberczki2020characteristic . Detailed dataset descriptions are provided in Appendix D.
Implementation Details. For T-Copilot, we construct the Copilot model using the same type of decoder architecture as the Pilot model to ensure consistency. We use the AdamW optimizer and Cosine learning rate scheduler for both Pilot and Copilot models. We modify the generate in HuggingFace Transformers huggingface-transformers to perform token-level logits fusion and rectified next-token generation during inference. All experiments are conducted on NVIDIA A100 GPUs. We leave all hyperparameter setups and training/inference details in Appendix E.1.
@@
Models and Baselines. We incorporate T-Copilot with varying backbone Pilot models. For encoder-decoder Pilots, we utilize T5 raffel2020exploring and FLAN-T5 chung2024scaling across small/base/large variants. For decoder-only Pilots, we employ multiple models from LLaMA-3 dubey2024llama and Qwen2.5 yang2024qwen2 families. We denote T-Copilot-small/base/0.5B/1B/3B as the Copilot model on different scales. Detailed model configuration and implementation details are provided in Appendix E.2. We compare against three baseline types: (i) Pilot-only models as described above. (ii) Frontier LLMs with comparable and larger parameters, including LLaMA-3.1-8B dubey2024llama , Gemma-2-9B team2024gemma , and Qwen2.5-14B. (iii) Layer/Adapter expansion methods, including MoE models shazeer2017outrageously (Mistral-7B, Ministral-8B), LLaMA/Mistral-Pro-8B wu2024llama , Mergekit-9B goddard2024arcee , and TIES yadav2024ties . Detailed baseline descriptions are provided in Appendix E.3.
@@ 5.1 Incorporating T-Copilot into Pilot Models Yields Better Performance
Comparison with Size-Matched Baselines. As shown in Table 2, we first compare our method against stronger models with larger parameters under the same model backbones. While LLaMA-3.2-3B initially lags significantly behind LLaMA-3.1-8B, incorporating T-Copilot-3B enables the model to outperform LLaMA-3.1-8B, despite using 2B fewer total parameters. Similarly, for the Qwen2.5 series, incorporating T-Copilot-3B enables the smaller Qwen2.5-7B to surpass Qwen2.5-14B with 4B fewer parameters. To provide a broader perspective, we also compare with strong baselines from different methods and model types. For instance, although LLaMA-3.2-3B originally trails behind models like Ministral-8B and LLaMA-Pro-8B, incorporating T-Copilot-3B enables it to outperform the strongest baseline under the 8B scale, Ministral-8B, while maintaining a 2B parameter advantage. Due to page limits, full comparison results are provided in Appendix F.1.
Downstream Tasks. Additional evaluation of T-Copilot and baseline comparisons on downstream recommendation tasks is provided in Appendix F.2.
@@ 5.2 Efficiency, Transferability, and Scalability
Efficiency. To thoroughly evaluate T-Copilot’s running efficiency, we compare against Pilot and baseline models with the same LLaMA-3 backbone architecture under similar parameter scales. As shown in Figure 5, T-Copilot maintains comparable inference throughput (Figure 5 (a)) and training speed (Figure 5 (b)) to its corresponding Pilot models, while incurring only a 4% marginal average increase in time overhead (Figure 5 (c)). In contrast, other baselines such as LLaMA-Pro-8B and MergeKit-9B suffer from significantly higher latency and computational costs relative to their base model LLaMA-3.1-8B. We provide a more detailed inference latency report in Appendix F.3 (Table 15) and discuss the architectural advantage of our model design in Appendix A.1.
T-Copilot-1B HellaSwag BoolQ GSM8K SVAMP Overall Impr.
with new LLaMA-3.2-1B 63.1 65.2 32.2 51.4 $\uparrow$ 6.1%
with new LLaMA-3.3-3B 91.4 70.2 58.8 68.5 $\uparrow$ 4.2%
with new LLaMA-3.1-8B 93.1 71.7 66.0 75.8 $\uparrow$ 2.4%
Transferability. In the T-Copilot learning framework, the Copilot model is fine-tuned alongside but separately from the Pilot model. Since the same type of models generally have similar learning trajectories under identical training settings, we further investigate if the Copilot model can leverage the mistake log of one Pilot model and still be effective on another Pilot model of the similar type.
We conduct controlled experiments on LLaMA-3 series models in which we directly apply a finetuned 1B Copilot model to new Pilot models during inference. The new Pilot model shares the same architecture as the original one but is trained independently. Note that the Copilot model does not "see" or "learn" any information from the new Pilot model, as they are not jointly trained during finetuning. In Table 3, transferring the Copilot model leads to a slight $\pm$0.2% performance difference compared to applying the Copilot to the initial Pilot models (jointly training together). We hypothesize that the minor discrepancy is due to the hardware inference differences between the original and new Pilot models. Nonetheless, the transferred Copilot model still delivers substantial performance gains for the new Pilot and consistently outperforms competing baselines. These results demonstrate that T-Copilot’s error-correction capabilities are not tightly coupled to a specific Pilot model and can be effectively transferred without additional rounds of fine-tuning.
Scalability. Figure 6 illustrates the relationship between accuracy and model parameter size for T-Copilot. Overall, incorporating the Copilot model consistently demonstrates improved performance as model size increases. We analyze the relationship between performance accuracy ($A$) and model parameter size ($N$) in billions. The derived equations for our method are as follows: for Flan-T5 backbones: $A\simeq 8.74\cdot\log_{10}(N)+40.17$; For LLaMA-3 backbones: $A\simeq 29.58\cdot\log_{10}(N)+50.20$, and for Qwen2.5 backbones: $A\simeq 12.40\cdot\log_{10}(N)+71.80$.
Ablation Studies. Detailed ablation studies on T-Copilot, including model design choices, input insertion patterns, and the effect of the hyperparameter $\lambda$, are presented in Appendix F.4.
@@ 6 Related Works
Recent research has explored various self-refinement techniques in LLMs to generate high-quality outputs. Models either iteratively prompt themselves with updated responses madaan2023self ; gulcehre2023reinforced ; shinn2023reflexion or optimize their behavior using external human or synthetic feedback ouyang2022training ; yao2023react ; miao2023selfcheck ; zou2025reasonflux . Orthogonal to external supervision such as additional prompting, multi-stage feedback, or explicit reward optimization, our work focuses on capturing model-internal signals during fine-tuning to achieve token-level rectification, without modifying the training objective or data distribution. We leave the additional related work and discussions in the Appendix G.
@@ 7 Conclusion
In this paper, we introduce Transformer Copilot, a novel learning framework that enhances Transformer-based Pilot models by integrating an auxiliary Copilot model during fine-tuning. By capturing the Pilot model’s learning signals in a Mistake Log during fine-tuning, the Copilot model learns to rectify the Pilot’s logits at inference time, enabling error-aware predictions. We provide both theoretical and empirical evidence that our method improves the Pilot model’s inference predictions. Experiments on 12 benchmarks demonstrate the effectiveness, efficiency, scalability, and transferability of Transformer Copilot. Discussions on limitations are provided in Appendix H.
Acknowledgment
We would like to thank the anonymous reviewers and AC for their constructive feedback during the review process. We also thank Sirui Chen, Tianxin Wei, Zhining Liu, Xiao Lin and other members from UIUC iDEA-iSAIL Lab for their constructive feedback on this work. This work is supported by National Science Foundation under Award No. IIS-2117902, and Agriculture and Food Research Initiative (AFRI) grant no. 2020-67021-32799/project accession no.1024178 from the USDA National Institute of Food and Agriculture. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.